# Merge then Realign: Simple and Effective Modality-Incremental Continual Learning for Multimodal LLMs

### v1

@@ 5.1 Experimental Setup
We build our MCL experiments on four modalities: image, video, audio, and point cloud, with two different training orders. Based on the prevalence of different modalities, we determine the two orders as follows. Sequential Order: image $\rightarrow$ video $\rightarrow$ audio $\rightarrow$ point cloud. Reverse Order: point cloud $\rightarrow$ audio $\rightarrow$ video $\rightarrow$ image. On top of this, the adopted datasets, metrics, models, and baselines are listed as follows.
Datasets. For each modality $M_{j}$, we leverage a dataset of Captioning (Cap) task and a dataset of Question Answering (QA) task to form the joint dataset $D_{j}=\{D_{j,Cap},D_{j,QA}\}$. The Cap and QA datasets for each modality are listed respectively. For image modality, we use MSCOCO-2014(Lin et al., 2014) and OK-VQA(Marino et al., 2019). For video modality, we use MSVD(Chen and Dolan, 2011) and MSVD-QA(Xu et al., 2017). For audio modality, we use AudioCaps(Kim et al., 2019) and Clotho-AQA(Lipping et al., 2022). For point cloud modality, we use a subset of Cap3D(Luo et al., 2024) and a subset of Cap3D-QA(Panagopoulou et al., 2023). More details of these datasets are in Appendix A.
Evaluation Metrics. First, we leverage Relative Gain(Scialom et al., 2022; Wang et al., 2024) as a normalized metric across different tasks. We train expert MLLMs individually on each single modality $M_{j}$ and test with their respective holdout data, taking their scores on the $k$-th dataset $D_{j,k}$ as upper bound $S_{j,k}^{sup}$. In the incremental stage $i$, the Relative Gain of modality $M_{i}$ with its dataset $D_{j}=\{D_{j,k}\}_{k=1}^{K}$ is calculated by:
@@
where $S_{j,k}^{i}$ is the score on the test set of $D_{j,k}$ in the stage $i$. Here, we utilize CIDEr score(Vedantam et al., 2015) and prediction accuracy (Acc) for Cap and QA tasks respectively to calculate $S_{j,k}^{sup}$ and $S_{j,k}^{i}$. To evaluate the performance degradations of the previously learned modalities, we calculate the Backward Relative Gain in the stage $i$ as:
@@
Model and Training Details. We leverage the mainstream MLLM architecture, i.e., LLaVA-like architecture with the Llama-3-8B-Instruct (Dubey et al., 2024) as its LLM backbone. The selections of modality encoders and connectors are detailed in Appendix B. Trainings that involve updating the LLM backbone utilize LoRA(Hu et al., 2022) for parameter-efficient fine-tuning. The training process in our merging stage is the same as the regular MLLM training, i.e., in the first step, only the connector is updated with Cap datasets, then in the second step, the connector and the LLM backbone are updated with all the task-related datasets (the combination of Cap and QA datasets in our case). In our realigning stage, the replay datasets are randomly sampled from the joint datasets of Cap and QA tasks. For each training process, the hyperparameters are listed in Appendix B.
@@
Figure 2: Progressive Backward Relative Gain in modality-incremental continual learning. For each stage $i$, we plot the average score of corresponding Backward Relative Gain with two different training orders. We set Backward Relative Gain to 100% for the $1$-st stage, denoting the initial performance without degradation.
@@
Baselines. In our experiments, we compare our MERA with non-CL fine-tuning and the representative methods of each CL category: Fine-Tuning: directly train MLLMs sequentially on each modality without applying any CL method. Replay: the vanilla replay-based CL method. During training on a new task, the model is updated with both samples from the current task and a set of randomly sampled replay data from previous tasks, to review knowledge of earlier tasks while learning the new one. EWC(Kirkpatrick et al., 2017): the most representative regularization-based CL method. EWC mitigates forgetting by restricting the updates of important weights during training on new tasks. It uses the Fisher information matrix to measure the importance of each weight. PathWeave(Yu et al., 2024a): an architecture-based CL method, also the first MCL method for MLLMs. PathWeave uses an adapter-in-adapter mechanism to memorize and extract knowledge from historical modalities to enhance the learning of the current modality. PathWeave is originally built on X-InstructBLIP(Panagopoulou et al., 2023). For a fair comparison, we implement PathWeave for our adopted MLLM architecture [^3] In their original settings, PathWeave removes the newly added modules when testing the former modalities. However, this results in the inability to perform cross-modality tasks, which are common in real applications. Therefore, we do not remove them for a fair comparison.. Implementation details of each baseline method are in Appendix D.
@@
Method Training Inference
Peak Memory Time-Consuming Peak Memory Lantency per Token
Fine-Tuning 37.43 GB 53 h 17.71 GB 34 ms
Replay (1%) 37.43 GB 54 h 17.71 GB 34 ms
Replay (10%) 37.43 GB 59 h 17.71 GB 34 ms
EWC 38.73 GB 54 h 17.71 GB 34 ms
PathWeave 40.08 GB 81 h 20.32 GB 111 ms
MERA (1%) 37.43 GB 54 h 17.71 GB 34 ms
MERA (10%) 37.43 GB 61 h 17.71 GB 34 ms
@@
Table 4: Ablation study of different components in MERA. The realigning stage uses 10% replay data. The mean and standard deviation of Backward Relative Gains in all the training stages are reported on different training orders. The best results are in bold.
5.2 Main Results
We conduct experiments under our MCL setting with both sequential and reverse orders. For Replay and our MERA, results using $r\%,r=\{1,10\}$ replay data are reported, denoted by Replay ($r\%$) and MERA ($r\%$) respectively. The progressive Backward Relative Gains averaged from different training orders are plotted in Figure 2. It is observed that our MERA demonstrates an impressive capability of mitigating performance degradation with consistent and promising Backward Relative Gains. When extending to all the four modalities, MERA ($10\%$) holds up to a 99.84% Backward Relative Gain, indicating that MERA can achieve a nearly lossless MCL performance, with at least 20.37% absolute improvements of Backward Relative Gain compared with other baselines. Notably, when only leveraging $1\%$ replay data, MERA (1%) can still achieve at least 15.94% absolute improvements over other baselines.
Further, we calculated the mean and standard deviation of Backward Relative Gains in all training stages for each method, in different training orders. Table 2 shows that our MERA (10%) achieves the highest mean and lowest standard deviation in both training orders, indicating its superior performance and high stability. Notably, in sequential order, MERA (10%) performs even better than lossless MCL, with an over 100% average Backward Relative Gain, also at least 14.15% absolute improvements of average Backward Relative Gain over other baselines. In reverse order, MERA (10%) also achieves at least 13.33% absolute improvements. When with only 1% replay data, MERA (1%) still achieves at least 11.05% and 4.33% absolute improvements in sequential and reverse orders respectively.
From Figure 2 and its raw data shown in Appendix E, we also observe a faint phenomenon of Positive Backward Transfer(Lin et al., 2022) that learning new knowledge improves the performance on previously learned tasks. For most CL methods and non-CL fine-tuning, the Backward Relative Gain comes to a low level when incrementally learning the first modality (corresponding to the $2$-nd training stage), but starts to stabilize or even increase when incrementally learning more modalities.
5.3 Efficiency Comparisons
We further compare the efficiency of different baselines and our MERA, as shown in Table 3. It is observed that our MERA can achieve optimal results except that MERA (1%) and MERA (10%) introduce 2% and 15% extra training time-consuming respectively. However, we believe its trade-off between training time-consuming and performance is worthwhile, considering the impressive performance of MERA. In our experiments, EWC and PathWeave introduce marginal extra training memory overhead, as we employ parameter-efficient fine-tuning. However, for larger LoRA ranks or even full model fine-tuning, their extra training memory consumptions would be substantial, as they necessitate storing additional parameters whose sizes increase linearly with the trainable parameters.
5.4 Ablation Study
We conduct ablation studies to investigate the effectiveness of each component in MERA. Results are shown in Table 4. Firstly, from Table 4 and Table 2, it is observed that the merging stage alone can already beat many other baselines, achieving the best and second-best performances among baselines in sequential and reverse orders respectively. Secondly, the realigning stage alone can also achieve performance improvements by addressing the misalignment issue. Combining both the merging and realigning stages, MERA further narrows the gap between the incrementally learned models and the individually trained experts on each modality, even, surpassing the individually trained experts in sequential training order with over 100% Backward Relative Gain.
@@ 5.4 Ablation Study
Table 5: Applying realigning to different CL or non-CL methods can further improve their Backward Relative Gain and stability. The realigning stage uses 10% replay data. The increased mean indicates better performance, while the decreased standard deviation indicates better stability. Improvements are colored in green.
Figure 3: Progressive Forward Relative Gain in modality-incremental continual learning. For each stage $i$, we plot the average score of corresponding Forward Relative Gain with two different training orders. We set Forward Relative Gain to 100% for the $1$-st stage, denoting the initial lossless plasticity.
5.5 Misalignment Is Common in MCL
Since the realigning stage achieves great success on top of our proposed merging stage, we further ask another question: does realigning benefit other CL methods, or is misalignment a common phenomenon in MCL? To examine this, we perform the realigning stage at the end of every training stage for different CL or non-CL methods to observe whether there are performance improvements. Table 5 shows that the additional realigning stage brings substantial performance improvements and increased stability for different CL or non-CL methods. Based on this observation, we can conclude that misalignment is a common phenomenon in MCL, and can be compensated by our proposed realigning stage.
5.6 Plasticity Analysis
Aside from alleviating performance degradation, the capability to adapt to new knowledge, i.e., plasticity, is also an important aspect. We use Forward Relative Gain as the metric to evaluate it. The progressive Forward Relative Gains averaged from different training orders are plotted in Figure 3. It is observed that the most elastic CL methods are EWC and PathWeave, while our MERA (10%) demonstrates comparable plasticity.
Notably, from Figure 3, we observe a strong phenomenon of Positive Forward Transfer(Ke et al., 2021) that the knowledge or skills acquired from earlier tasks improve the learning efficiency of new tasks. The Positive Forward Transfer emerges before the $4$-th training stage, on EWC, PathWeave, and MERA. This phenomenon is also reported by other MCL literature(Yu et al., 2024a). In contrast to Positive Forward Transfer, there is a gradual plasticity loss(Dohare et al., 2024, 2023) as the model attempts to retain more knowledge. This explains the decreases in Forward Relative Gain across different CL methods in the $4$-th stage, as the plasticity loss comes to a dominant position.
@@ 6 Future Work
This work is one of the early attempts of MCL, mainly focusing on designing effective methods and addressing the misalignment issue. However, our extensive experiments suggest that there are many other interesting aspects of MCL that warrant attention. In Section 5.2 and Section 5.6, we observe the phenomenons of Positive Backward Transfer and Positive Forward Transfer, respectively. This could be related to the complex cross-modal interaction in multimodal learning, urging for further research on the mechanisms of modality interaction and methods that boost a positive modality interaction, in the context of MCL.
@@ 7 Conclusion
In this paper, we mainly focus on designing effective methods for Modality-incremental Continual Learning (MCL). First, we revisit MCL and investigate a more severe issue it faces in contrast to traditional continual learning that its performance degradation comes not only from forgetting but also from misalignment. To address both the forgetting and misalignment, we propose MERA, a simple yet effective MCL paradigm. Extensive experiments demonstrate that MERA significantly outperforms the baselines, and even achieves a nearly lossless MCL performance. Further, we observe a sign of complex cross-modal interaction in MCL, providing a direction for future work.
Limitations
Our work is restricted in the following aspects. First, our experiments are limited to four commonly used modalities due to the lack of resources for other less-studied modalities. Second, our experiments are limited to two types of tasks for each modality due to the lack of general-purpose and diverse instruction datasets and their corresponding benchmarks for some modalities, i.e., audio and point cloud. Third, our method is limited to any-to-text MLLMs while there is now a trend of exploring any-to-any MLLMs. However, the main idea of MERA is generic to any-to-any MLLMs since their architecture topologies are similar. Fourth, our work only explores the simplest model merging method in the context of MCL, aiming to provide a universal framework, leaving the adaptation of other model merging methods to MCL for future work.
@@
Appendix A Dataset Details
Appendix B Implementation Details
[…]
  • Point Cloud: We use Point-BERT-v1.2Xu et al. (2024) as the pre-trained point cloud encoder, a randomly initialized MLP as the connector.
We set the hyperparameters mainly following previous worksLiu et al., 2024b, a, as listed in Table 7. For training that involves updating the LLM backbone, we utilize parameter-efficient fine-tuning with LoRA(Hu et al., 2022) applied across all linear modules within the LLM, setting the LoRA rank to 128 and the alpha parameter to 128. All the experiments are conducted on a single NVIDIA RTX A6000 48G with FP16.

### v2

@@ 5.1 Experimental Setup
We build our MCL experiments on four modalities: image, video, audio, and point cloud, with two different training orders. Based on the prevalence of different modalities, we determine the two orders as follows. Sequential Order: image $\rightarrow$ video $\rightarrow$ audio $\rightarrow$ point cloud. Reverse Order: point cloud $\rightarrow$ audio $\rightarrow$ video $\rightarrow$ image. On top of this, the adopted datasets, metrics, models, and baselines are detailed as follows.
Datasets. For each modality $M_{j}$, we leverage a dataset of Captioning (Cap) task and a dataset of Question Answering (QA) task to form the joint dataset $D_{j}=\{D_{j,Cap},D_{j,QA}\}$. The Cap and QA datasets for each modality are listed respectively. For image modality, we use MSCOCO-2014 (Lin et al., 2014) and OK-VQA (Marino et al., 2019). For video modality, we use MSVD (Chen and Dolan, 2011) and MSVD-QA (Xu et al., 2017). For audio modality, we use AudioCaps (Kim et al., 2019) and Clotho-AQA (Lipping et al., 2022). For point cloud modality, we use a subset of Cap3D (Luo et al., 2024) and a subset of Cap3D-QA (Panagopoulou et al., 2023). More details of these datasets are in Appendix B.
Evaluation Metrics. First, we leverage Relative Gain (Scialom et al., 2022; Wang et al., 2024b) as a normalized metric across different tasks. We naively train (without using any continual learning methods) expert MLLMs individually on each single modality $M_{j}$ and test with their respective holdout data, taking their scores on the $k$-th dataset $D_{j,k}$ as upper bound $S_{j,k}^{sup}$. In the incremental stage $i$, the Relative Gain of modality $M_{i}$ with its dataset $D_{j}=\{D_{j,k}\}_{k=1}^{K}$ is calculated by:
@@
where $S_{j,k}^{i}$ is the score on the test set of $D_{j,k}$ in the stage $i$. Here, we utilize CIDEr score (Vedantam et al., 2015) and prediction accuracy (Acc) for Cap and QA tasks respectively to calculate $S_{j,k}^{sup}$ and $S_{j,k}^{i}$. To evaluate the performance degradation of the previously learned modalities, we calculate the Backward Relative Gain in the stage $i$ as:
@@
Figure 3: Progressive Backward Relative Gain in modality-incremental continual learning. For each stage $i$, we plot the average score of the corresponding Backward Relative Gain with two different training orders. We set Backward Relative Gain to 100% for the 11st stage, denoting the initial performance without degradation. Exceptionally, the initial Backward Relative Gain of EProj is not 100% since it only tunes the modality-specific components, causing an initial performance degradation.
Model and Training Details. We leverage the mainstream MLLM architecture, i.e., LLaVA-like architecture with the Llama-3-8B-Instruct (Dubey et al., 2024) as its LLM backbone. The selections of modality encoders and connectors are detailed in Appendix C.1. Trainings that involve updating the LLM backbone utilize LoRA (Hu et al., 2022) for parameter-efficient fine-tuning. The training process in our merging stage is the same as the standard MLLM training, i.e., in the first step, only the connector is updated with Cap datasets, then in the second step, the connector and the LLM backbone are updated with all the task-related datasets (the combination of Cap and QA datasets in our case). In our realigning stage, the replay datasets are randomly sampled from the joint datasets of Cap and QA tasks. For each training process, the hyperparameters are listed in Appendix C.1.
@@
Baselines. In our experiments, we compare our MERA with non-CL fine-tuning, as well as the representative CL and MCL methods: Fine-Tuning: directly train MLLMs sequentially on each modality without applying any CL method. Replay: the vanilla replay-based CL method. During training on a new task, the model is updated with both samples from the current task and a set of randomly sampled replay data from previous tasks. EWC (Kirkpatrick et al., 2017): the most representative regularization-based CL method. EWC mitigates forgetting by restricting the updates of important weights during training on new tasks. It uses the Fisher information matrix to measure the importance of each weight. EProj (He et al., 2023): tuning only the modality-specific components to prevent forgetting. PathWeave (Yu et al., 2024a): an architecture-based CL method, also the state-of-the-art MCL method for MLLMs. PathWeave uses an adapter-in-adapter mechanism to memorize and extract knowledge from historical modalities to enhance the learning of the current modality. PathWeave is originally built on X-InstructBLIP (Panagopoulou et al., 2023). For a fair comparison, we implement PathWeave for our adopted MLLM architecture. Implementation details of each baseline method are in Appendix C.3.
Figure 4: Progressive Forward Relative Gain in modality-incremental continual learning. For each stage $i$, we plot the average score of the corresponding Forward Relative Gain with two different training orders. We set Forward Relative Gain to 100% for the 11st stage, denoting the initial lossless plasticity. Exceptionally, the initial Forward Relative Gain of EProj is not 100% since it only tunes the modality-specific components, causing an initial loss of plasticity.
5.2 Main Results
We conduct experiments under our MCL setting with both sequential and reverse orders. For Replay and our MERA, results using rr%, $r=\{1,10\}$ replay data are reported, denoted by Replay (rr%) and MERA (rr%) respectively.
Evaluating Degradation. The progressive Backward Relative Gains averaged from different training orders are plotted in Figure 3. It is observed that our MERA demonstrates an impressive capability of mitigating performance degradation with consistent and promising Backward Relative Gains. When extending to all four modalities, MERA ($10\%$) holds up to a 99.84% Backward Relative Gain, indicating that MERA can achieve nearly lossless MCL performance, with at least 12.09% absolute improvements of Backward Relative Gain compared with other baselines. Notably, when only leveraging $1\%$ replay data, MERA (1%) can still achieve at least 7.66% absolute improvements over other baselines. Further, we calculated the mean and standard deviation of Backward Relative Gains in all training stages for each method, in different training orders. Table 2 shows that our MERA (10%) achieves the highest mean in both training orders, indicating its superior performance. It also achieves the lowest and second-lowest standard deviation in sequential and reverse orders respectively, indicating its high stability. Notably, in sequential order, MERA (10%) performs even better than lossless MCL, with an over 100% average Backward Relative Gain, also at least 11.39% absolute improvements of average Backward Relative Gain over other baselines. In reverse order, MERA (10%) also achieves at least 13.33% absolute improvements. When with only 1% replay data, MERA (1%) still achieves at least 8.29% and 4.33% absolute improvements in sequential and reverse orders respectively.
@@ 5.2 Main Results
Method Sequential Reverse
Fine-Tuning 38.50 36.14
Replay (1%) 30.61 33.02
Replay (10%) 56.68 32.50
EWC 25.79 24.94
EProj 55.61 55.61
PathWeave 36.32 46.39
MERA (1%) 72.13 61.22
MERA (10%) 72.70 54.04
Table 3: Accuracies on MCUB-4 benchmark. Results are reported on different training orders and are measured after continually learning on all four modalities.
Evaluating Degradation on Complex Multimodal Tasks. To evaluate MERA’s performance in more complex multimodal tasks that involve multiple modalities at a time, we further conduct experiments on the MCUB-4 (Chen et al., 2024) benchmark that requires the model to simultaneously infer on all four modalities, i.e., image, video, audio, and pointcloud. Table 3 shows the accuracies of each method on MCUB-4 benchmark. It is observed that our proposed MERA also significantly excels in more complex multimodal tasks that involve multiple modalities at a time.
Evaluating Plasticity. Aside from alleviating performance degradation, the capability to adapt to new knowledge, i.e., plasticity, is also an important aspect. We use the Forward Relative Gain as the metric. The progressive Forward Relative Gains averaged from different training orders are plotted in Figure 4. It is observed that the most elastic CL methods are EWC and PathWeave, while our MERA (10%) demonstrates comparable plasticity.
5.3 Ablation Study
We conduct ablation studies to investigate the effectiveness of each stage of MERA. Results are shown in Table 4. Firstly, from Table 4 and Table 2, it is observed that the realigning stage that addresses the misalignment issue can already beat many other baselines designed to tackle the forgetting issue, achieving the second-best average Backward Relative Gain among baselines in both sequential and reverse orders. Secondly, combining both merging and realigning stages, MERA further narrows the gap between the incrementally learned models and the individually trained experts on each modality, even, surpassing the individually trained experts in sequential training order with over 100% Backward Relative Gain.
@@ 5.3 Ablation Study
Table 4: Ablation study of different components in MERA. The realigning stage uses 10% replay data.
6 Discussions
6.1 Is Misalignment Common in MCL?
Since the realigning stage achieves great success on top of our proposed merging stage, we further ask another question: does realigning benefit other CL methods, or in other words, is misalignment a common phenomenon in MCL? To examine this, we perform the realigning stage at the end of every training stage for different CL or non-CL methods to observe whether there are performance improvements [^3] We do not examine this for EProj. EProj is self-evidently misalignment-free, since it freezes the LLM backbone.. Table 5 shows that the additional realigning stage brings substantial performance improvements and increased stability for different CL or non-CL methods. Based on this observation, we can conclude that misalignment is a common phenomenon in MCL, and can be alleviated by our proposed realigning stage.
@@ 6.1 Is Misalignment Common in MCL?
Table 5: Applying realigning to different CL methods can further improve their Backward Relative Gain and stability. The realigning stage uses 10% replay data. Improvements are colored in green.
6.2 Positive Backward Transfer and Positive Forward Transfer
From Figure 3 and its raw data shown in Appendix D, we observe a faint phenomenon of Positive Backward Transfer (Lin et al., 2022) that learning new knowledge improves the performance on previously learned tasks. For most methods, the Backward Relative Gain comes to a low level when incrementally learning the second modality, but starts to stabilize and even increase when incrementally learning more modalities.
From Figure 4, we observe a strong phenomenon of Positive Forward Transfer (Ke et al., 2021) that the knowledge acquired from earlier tasks improves the learning efficiency of new tasks. The Positive Forward Transfer exists before the $4$-th incremental stage, when employing EWC, PathWeave, and MERA. This phenomenon is also reported in other MCL literature (Yu et al., 2024a). In contrast to Positive Forward Transfer, there is a gradual loss of plasticity (Dohare et al., 2024, 2023) as the model attempts to retain more knowledge. This explains the decreases in Forward Relative Gain across different CL methods in the $4$-th stage, as the loss of plasticity comes to a dominant position.
@@ 7 Future Work
This work is one of the early attempts of MCL for MLLMs, focusing on the dual causes of its degradation. In addition to this, observation of Positive Backward Transfer and Positive Forward Transfer in Section 6.2 may imply the complex cross-modal interaction in multimodal learning, urging for future research on the mechanisms of modality interaction in the context of MCL.
@@ 8 Conclusion
In this paper, we first revisit MCL and uncover the dual causes of its degradation, i.e., forgetting and misalignment. Next, to address both forgetting and misalignment, we propose MERA, a simple yet effective MCL paradigm. Extensive experiments demonstrate that MERA significantly outperforms the state-of-the-art methods and even achieves nearly lossless MCL performance. Our findings underscore the misalignment issue in MCL. More broadly, our work showcases how to adjust different components of MLLMs during continual learning. Further, we observe signs of complex cross-modal interaction in MCL, providing a direction for future work.
9 Limitations
This work is restricted in the following aspects. First, our experiments are limited to four commonly used modalities due to the lack of resources for other less-studied modalities. Second, we limit this work to LLaVA-like architecture as it covers the majority of MLLMs. Third, this work is limited to any-to-text MLLMs while there is now a trend of exploring any-to-any MLLMs. However, the main idea of MERA is generic to them. Fourth, our work only explores the simplest model merging method in the context of MCL, aiming to provide a universal framework, leaving other model merging methods for MCL for future work.
10 Acknowledgments
This work was supported by the National Natural Science Foundation of China (No.62372139), the National Natural Science Foundation of China (2024A1515030024), Research Projects of Shenzhen (JCYJ20220818102414030), and Key Laboratory of Guangdong Province (2022B1212010005).
@@ 10 Acknowledgments
Appendix A Further Analysis and Discussions
A.1 Deeper Discussions on the Distinction Between Forgetting and Misalignment
Theoretically, forgetting is associated with various factors such as representation drift (Caccia et al., ), gradient interference (Wang et al., 2023), learning dynamics (Ren and Sutherland, 2024), distribution shift (which causes misalignment), making it a comprehensive issue. For example, Schwettmann et al. (2023) finds that there exist multimodal neurons in the LLM backbone, each corresponds to certain domain concepts, and when breaking some of these neurons, the model’s prediction is largely affected. However, addressing the misalignment issue by our proposed realigning will possibly not recover the model’s performance since the connectors do not encode domain-specific concepts (Verma et al., 2024). This example showcases that the forgetting issue is not limited to misalignment. Empirically, our experiments in Table 5 have shown that, a) addressing only the misalignment issue, i.e., fine-tuning with realigning, does not outperform some methods that address the overall forgetting, e.g., PathWeave/merging without realigning, b) methods that address the overall forgetting significantly complement the realigning alone (fine-tuning with realigning). In terms of methodology, forgetting is a comprehensive issue that can be tackled mainly by heuristic methods such as EWC or replay, while misalignment is caused solely by distribution shift that can be directly addressed by realigning the LLM and modality-specific components. In terms of methodology, forgetting is a comprehensive issue that can be tackled mainly by heuristic methods such as EWC or replay, while misalignment is caused solely by distribution shift that can be directly addressed by realigning the LLM and modality-specific components.
A.2 Misalignment Also Exists In Replay-Based Methods
Theoretically, a potential exception to misalignment is when replay-based methods are applied, since the $\theta^{LLM}$ is trained on the joint distribution of all $\{M_{1},M_{2},\dots,M_{i}\}$ modalities. However, Table 5 suggests that replay-based methods still suffer from misalignment and can be compensated for by our proposed realigning stage. We conjecture that it is due to the imbalanced distribution of its training data. Replay-based methods train the model on the joint dataset of its replay data and $D_{i}$, where the scale of replay data from each previous modality is insignificant to the scale of $D_{i}$. Therefore, it still suffers from a certain degree of distribution shift during its training process.
A.3 Easily Adapt MERA to Other MLLM Architectures
Although we limit our work to LLaVA-like architecture, our method can be easily adapted to other MLLM architectures. We present several naive ways to adapt MERA to some other MLLM architectures.
  • For connector-free MLLMs, the last few layers of their encoders can be treated as connectors so that our method can be directly applied.
  • For MLLMs that use a uni-connector for all modalities, we can treat them as the connector-free MLLMs.
  • For encoder-free MLLMs, we can realign the raw multimodal input distributions instead of multimodal feature distributions with the LLM backbone by fine-tuning the embedding layers.
A.4 Efficiency Comparisons
We compare the efficiency of different baselines and our MERA, as shown in Table 6. It is observed that our MERA can achieve optimal results except that MERA (1%) and MERA (10%) introduce 2% and 15% extra training time-consuming respectively. However, we believe its trade-off between training time-consuming and performance is worthwhile, considering the impressive performance of MERA. In our experiments, EWC and PathWeave introduce marginal extra training memory overhead, as we employ parameter-efficient fine-tuning. However, for larger LoRA ranks or even full model fine-tuning, their extra training memory consumptions would be substantial, as they necessitate storing additional parameters whose sizes increase linearly with the trainable parameters.
Method Training Inference
Peak Memory Time-Consuming Peak Memory Lantency per Token
Fine-Tuning 37.43 GB 53 h 17.71 GB 34 ms
Replay (1%) 37.43 GB 54 h 17.71 GB 34 ms
Replay (10%) 37.43 GB 59 h 17.71 GB 34 ms
EWC 38.73 GB 54 h 17.71 GB 34 ms
PathWeave 40.08 GB 81 h 20.32 GB 111 ms
MERA (1%) 37.43 GB 54 h 17.71 GB 34 ms
MERA (10%) 37.43 GB 61 h 17.71 GB 34 ms
Method Sequential Reverse
Mean Std Mean Std
MERA 97.90 6.02 84.42 12.93
MERA
w/ 10% noise
96.87 4.97 83.56 13.48
MERA
w/ 50% noise
94.10 5.67 79.79 15.54
Table 7: Evaluations of MERA’s robustness to noisy replay data. The realigning stage uses 10% replay data.
A.5 Robustness to the Quality of Replay Data
Since the realigning stage of MERA relies on a small replay dataset, it is beneficial to understand how robust MERA is to the quality of replay data. To evaluate its robustness, we manually corrupt $p\%$ samples in the replay dataset by mispairing their text-modality pair, and test MERA’s performance. Results in Table 7 suggest that 10% noisy samples does not significantly corrupt MERA’s performance, and it still shows decent performance even under the extreme condition of 50% noisy samples. This suggests that our MERA is robust to low-quality replay data.
A.6 Why is the Performance Degradation More Severe in Reverse Training Order
In Table 2, it is observed that the performance degradation is more severe in reverse training order than in sequential training order. This phenomenon occurs across different methods. We conjecture that training in sequential order is a sort of curriculum learning, while the reverse order corresponds to reversed curriculum learning. In our experiments, the sequential order is determined based on the prevalence of different modalities, and more prevalent modalities might be easier to learn. For example, the image modality is easier to understand than video, and the point cloud might be the most difficult modality to learn. Wang et al. (2024b) also observed that the performance degradation is less severe when continually learning in a curriculum learning order than in a reversed curriculum learning order. We conjecture that learning easy knowledge leads to the forgetting of harder ones, but learning hard knowledge might even consolidate the easier ones.